### v1

# RealCamo: Boosting Real Camouflage Synthesis with Layout Controls and Textual-Visual Guidance

Abstract
Camouflaged image generation (CIG) has recently emerged as an efficient alternative for acquiring high-quality training data for camouflaged object detection (COD). However, existing CIG methods still suffer from a substantial gap to real camouflaged imagery: generated images either lack sufficient camouflage due to weak visual similarity, or exhibit cluttered backgrounds that are semantically inconsistent with foreground targets. To address these limitations, we propose ReamCamo, a unified out-painting based framework for realistic camouflaged image generation. ReamCamo explicitly introduces additional layout controls to regulate global image structure, thereby improving semantic coherence between foreground objects and generated backgrounds. Moreover, we construct a multi-modal textual–visual condition by combining a unified fine-grained textual task description with texture-oriented background retrieval, which jointly guides the generation process to enhance visual fidelity and realism. To quantitatively assess camouflage quality, we further introduce a background–foreground distribution divergence metric that measures the effectiveness of camouflage in generated images. Extensive experiments and visualizations demonstrate the effectiveness of our proposed framework.
Code and results will be released soon.
Camouflaged Image Generation, Diffusion Model, Controlled Image Generation, Synthesize Data
Figure 1: Motivation and qualitative comparison of camouflaged image synthesis. (a) LAKE-RED (CVPR2024LAKERED) fails to achieve effective camouflage, with targets remaining visually distinguishable, whereas our method produces images with well-camouflaged targets. (b) CamoAny (CVPR2025CamoAny) achieves visual camouflage but suffers from weak semantic consistency between the target and the generated background, in contrast, our method preserves semantic coherence, resulting in more realistic scenes. (c) Our approach is capable of generating diverse and realistic camouflaged images while jointly satisfying visual similarity and semantic consistency.
1 Introduction
Why Synthesize Data: Concealed visual perception tasks, exemplified by camouflaged object detection (COD) (ICML2025RUN), aim to identify semantically meaningful targets that are visually indistinguishable from their surrounding backgrounds. Such tasks are increasingly important in real-world applications, including medical diagnosis, agricultural and industrial production (AIR2024CODsurvey). Despite recent progress, the development of COD remains fundamentally constrained by the limited scale of available training data.
In contrast to large-scale benchmarks for general object detection & segmentation, such as PASCAL VOC (IJCV2010PASCAL) and MS-COCO (ECCV2014MScoco), which contain tens to hundreds of thousands of precisely annotated images, existing COD datasets (CVPR2020SINet-COD10K; CVPR2021LSR-NC4K) typically comprise only a few thousand samples. This scarcity arises from two intrinsic challenges: (i) camouflaged objects are inherently difficult to annotate due to its high visual similarity to backgrounds, and (ii) camouflaged scenes with high quality are themselves rare and costly to capture in the real world. As a result, data insufficiency has become a critical bottleneck for improving both the robustness and generalization of COD models.
Recent advances in image generation models, including GAN (ACM2020GAN) and diffusion models (NIPS2020DDPM; CVPR2022LDM), offer a promising alternative, while leveraging high-quality synthetic data to augment training has been shown to be an effective and cost-efficient strategy in multiple vision tasks such as salient object detection (arXiv2025S3OD). This observation motivates the camouflaged image synthesis task, whose core objective is to generate realistic and diverse camouflaged scenes that can meaningfully complement limited real-world datasets and thereby improve downstream COD performance.
Figure 2: The proposed KL divergence based metric. Existing generative evaluation metrics, such as FID and KID, are insufficient to measure the similarity between foreground and background, while our proposed $\mathrm{KL}_{BF}$ provides a quantitative assessment of camouflage. Top: Image with a higher $\mathrm{KL}_{BF}$. Bottom: Our generated image with a lower $\mathrm{KL}_{BF}$, more camouflaged.
Why Out-Painting: Existing approaches to camouflaged image synthesis can be broadly categorized into two paradigms. The first paradigm, exemplified by style-transfer based methods such as LCGNet (TMM2022LCGNet), modifies the appearance of a target object to match a given background. However, this process often alters object-specific textures and visual cues to such an extent that the semantic identity of the target is partially or entirely compromised. Consequently, these methods achieve camouflage by enforcing visual uniformity, rather than by modeling the natural interaction between an object and its environment.
The second paradigm, represented by out-painting based methods such as LAKE-RED (CVPR2024LAKERED), follows a fundamentally different strategy. Instead of altering the target object, these methods preserve its semantic content and generate surrounding background regions whose textures and structures are consistent with the target appearance. This formulation enables camouflage to emerge from contextual compatibility between the object and its environment, rather than from direct manipulation of the object itself.
Compared with style-transfer based approaches, out-painting offers several advantages. First, it maintains the semantic integrity of the target while avoiding the need for an externally provided background image. More importantly, it aligns more closely with the natural principles of camouflage observed in the real world. Camouflage is inherently a relative and environment-dependent phenomenon: the same object may appear well camouflaged in one context but highly salient in another (arXiv2025COD-SOD). From the perspective of natural selection, creatures do not actively redesign their appearance to match each environment, instead, environmental variation favors the survival of individuals whose appearances are already compatible with their surroundings. This perspective naturally corresponds to the out-painting formulation.
Why Background Control: For out-painting based methods to produce realistic and effective camouflaged training data, we argue that two key requirements must be simultaneously satisfied. First, the generated background should exhibit strong appearance compatibility with the target, particularly in terms of color distribution and texture patterns, so that the target is genuinely camouflaged under visual perception. Second, the background must convey semantically plausible content consistent with real-world priors, for instance, aquatic objects should appear in water rather than in visually incompatible contexts.
Meeting these requirements is inherently challenging because camouflaged targets typically occupy only a small spatial region of the image. Consequently, out-painting methods must synthesize large background areas with limited target-conditioned information, often leading to a trade-off between visual similarity and semantic consistency. As shown in Fig. 1 (a & b), existing approaches exhibit clear deficiencies: some generated backgrounds that are insufficiently compatible with the target appearance, resulting in weak camouflage, while others produce visually plausible textures that lack coherent or meaningful semantics.
These limitations indicate that unconstrained out-painting is inadequate for high-fidelity camouflaged image generation. To address this issue, explicit guidance and control mechanisms during background synthesis are essential, enabling the generation process to balance appearance-level camouflage with semantic realism.
Our Focus and Contributions: Based on the above analysis, this work focuses on three tightly coupled contributions:
(1) Controlled camouflaged image synthesis. To overcome the limitations of existing out-painting based methods in background generation, we introduce a unified out-painting based framework termed RealCamo, which integrates explicit layout controls and textual-visual guidance. We argue that background layout plays a critical role in determining semantic plausibility. Building upon the contrast control in CamoAny, we incorporate additional depth and holistically-nested edge information into ControlNet to explicitly constrain the spatial layout of generated scenes, thereby preserving semantic realism. Furthermore, to enhance camouflage effectiveness at the texture level, we propose a textual–visual guidance mechanism that combines a unified fine-grained textual task descriptions with additional visual cues retrieved according to target texture characteristics, enhancing visual fidelity and realism of generation.
(2) Camouflage effectiveness evaluation metrics. A realistic camouflaged image should preserve the original scene layout while achieving camouflage through modifications in background color and texture. Existing generative evaluation metrics, such as FID and KID, are insufficient to capture this requirement. We therefore introduce SSIM to quantify structural consistency between generated and original images as a indicator for semantic realism. In addition, we propose a KL divergence based metric $\mathrm{KL}_{BF}$, which measures the similarity between background and foreground pixel distributions, providing a principled quantitative assessment of camouflage effectiveness, as shown in Fig. 2.
(3) Downstream evaluation. To assess the practical utility of the generated camouflaged images, we construct a large-scale synthetic dataset, SynCOD12K, as supplementary training data for camouflaged object detection. Extensive experiments on four COD benchmarks demonstrate consistent performance improvements, with further investigations on the effects of camouflage-augmented data on related salient and general object detection tasks.
2 Related Work
Synthetic Dataset Generation. Synthetic data has become an effective strategy for alleviating data scarcity in computer vision (MIR2024Synsurvey). Advances in generative models, including GANs and diffusion models, have enabled the construction of large-scale, high-quality synthetic datasets. DatasetGAN (CVPR2021DatasetGAN) generates image-annotation pairs by exploiting the latent feature space of pretrained GANs, while BigDatasetGAN (CVPR2022BigDatasetGAN) extends this paradigm to ImageNet-scale class diversity using VQGAN (CVPR2021VQGAN). More recently, diffusion-based methods such as DiffuMask (ICCV2023DiffuMask) and DatasetDM (CVPR2022LDM) leverage text-to-image models to produce photorealistic images and corresponding annotations for downstream perception tasks. Despite their success, these approaches primarily focus on generic image synthesis and annotation, and remain limited when applied to structurally complex and semantically subtle scenarios such as camouflaged image generation.
Camouflaged Image Generation. Camouflaged image generation, introduced by TOG2010Chu, can be categorized into two paradigms: Style transfer based background-guided approaches modify foreground appearance to blend targets into given backgrounds. DCI (AAAI2020DCI) employs an attention-aware camouflage loss to suppress salient cues while preserving subtle perceptual features, and LCGNet (TMM2022LCGNet) achieves extreme camouflage by fusing high-level foreground and background representations. In contrast, out-painting based foreground-guided methods generate backgrounds conditioned on the foreground target. LAKE-RED (CVPR2024LAKERED) introduces a knowledge retrieval–augmented framework using a pretrained codebook, while FACIG (ICME2025FACIG) refines feature integration to reduce foreground distortion CamoAny (CVPR2025CamoAny) further improves background synthesis through controlled out-painting and representation engineering. Despite these advances, existing methods still struggle to simultaneously achieve strong visual similarity and semantic consistency. More recently, the text-driven method CT-CIG (AAAI2026CTCIG) has shown promise in realistic camouflage synthesis, but the inability to provide aligned annotations limits their applicability to downstream camouflaged object detection applications.
Controlled Image Generation. Recent advances in text-to-image (T2I) generation integrate large-scale text encoders, such as CLIP (ICML2021CLIP) and T5 (JMLR2020T5), into diffusion models, yielding powerful T2I models. To enable fine-grained spatial control, ControlNet (ICCV2023ControlNet) introduces an auxiliary trainable branch to inject conditional signals into pretrained diffusion models. Building upon this design, UniControl (NIPS2023UniControl) employs a task-aware hypernetwork to modulate zero convolutions of the ControlNet in handling multiple conditions, while CtrLoRA (ICLR2025CtrLoRA) further reduces parameter overhead through condition-specific LoRA (ICLR2022LoRA) fine-tuning, and ICCV2025PreserveAnything proposes an N-channel ControlNet to improve object preservation, prompt alignment, and aesthetic quality under multiple control settings.
3 Methodology
3.1 Preliminary
Our proposed method is based on the classic Latent Diffusion Models (LDMs) (CVPR2022LDM) and ControlNet (ICCV2023ControlNet). Here, we briefly revisiting their basics, with more details can be found in Appendix B.
Figure 3: The pipeline of our camouflaged image generation framework RealCamo. Explicit layout controls (i.e., the contrast, depth, and HED controls) are extracted by the Layout Controls Generator (LCG) and injected to ControlNet, while the multi-modal condition produced by the Textual-Visual Condition Generator (TVCG) is used as a guidance in refining the camouflage generation process.
Latent Diffusion Models. LDMs generate high-quality images by reversing a Markov forward process in the latent space. Specifically, Given a VAE (ICML2014VAE2) with encoder $\mathcal{E}$ and decoder $\mathcal{D}$, Gaussian noise $\epsilon$ is progressively added to a latent representation $z_{0}=\mathcal{E}(x_{0})$ of the original image $x_{0}$, resulting in a noisy image $z_{t}=\alpha_{t}z_{0}+(1-\alpha_{t})\epsilon$, with more noise being introduced when $\alpha_{t}$ decreases as timestep $t$ increases. A denoising network $\epsilon_{\theta}$ is trained to predict the added noise given a condition $c^{t}$ following:
$\mathcal{L}=\mathbb{E}_{x_{0},t,\epsilon\sim\mathcal{N}(0,I),\mathcal{E}}\left[\parallel\epsilon-\epsilon_{\theta}(z_{t},t,c^{t})\parallel\right],$ (1)
and reverse the noise image to a latent representation $z^{\prime}_{0}$, which is finally decoded into the image space by $\mathcal{D}(z^{\prime}_{0})$.
ControlNet. ControlNet extends the structural controllability of LDMs by incorporating addition conditions ($c^{f}$) like Canny edges and human pose. With the introduction of additional trained branch aided by zero convolutions, ControlNet reformulates Equation 1 as:
$\mathcal{L}=\mathbb{E}_{x_{0},t,c^{t},c^{f},\epsilon\sim\mathcal{N}(0.I),\mathcal{E}}\left[\parallel\epsilon-\epsilon_{\theta}(z_{t},t,c^{t},c^{f})\parallel_{2}^{2}\right],$ (2)
which enables fine-grained spatial control during generation.
3.2 Framework Overview
As illustrated in Fig. 3, the proposed RealCamo framework is built upon Stable Diffusion 1.5 (CVPR2022LDM) and ControlNet (ICCV2023ControlNet), enabling controllable camouflaged image generation through complementary structural and semantic conditioning. Control signals are injected in two ways: (i) structural constraints introduced via ControlNet to regulate spatial layout, and (ii) multi-modal prompts injected through cross-attention to guide image content. These mechanisms align with our objective that high-quality camouflage requires both visual similarity and semantic consistency between foreground targets and background scenes. Accordingly, we employ a Layout Controls Generator to provide explicit structural priors via ControlNet, and a Textual–Visual Condition Generator that integrates a unified fine-grained task description with background images retrieved by a Texture-Oriented Background Retrieval module, producing informative multi-modal prompts to effectively steer the camouflage generation process.
3.3 Layout Controls Generator
We argue that semantic consistency in camouflaged image generation is largely determined by the structural layout of the image. Prior methods (CVPR2024LAKERED; CVPR2025CamoAny) often yield cluttered or physically implausible backgrounds due to the lack of explicit layout constraints. To address this, we propose a layout conditions generator (LCG) that extracts complementary structural cues and injects them via ControlNet, as shown in Fig. 3. Specifically, for a given image $\mathrm{I}$ and a corresponding mask $\mathrm{M}$ which indicate the foreground target, we employ LCG to extract contrast, depth, and holistically-nested edge (HED) controls as explicit structural priors to promote coherent spatial organization and improved foreground–background semantic alignment:
$c_{cst},c_{dep},c_{hed}=LCG(\mathrm{I},\mathrm{M},is\_training).$ (3)
Each control signal is embedded using a dedicated MLP (NIPS2023UniControl) due to distributional differences, and then fused as a structural guidance for ControlNet to process.
Contrast Control. Following CVPR2025CamoAny, contrast control $c_{cst}$ aims to preserve the foreground target while providing basic regulation of color contrast, varies when training and inference based on is_trainingis\_training.
Depth Control. Depth control $c_{dep}$ is generated using Depth Anything V2 (NIPS2024DepthAnythingV2), and is intended to constrain the spatial relationships within the image.
HED Control. HED control $c_{hed}$ is obtained by deploying Holistically-nested edge detection (ICCV2015HED) to regulate the contours of the main elements in the image.
3.4 Texture-Oriented Background Retrieval
Compared to generic image synthesis, camouflaged image generation requires backgrounds that are quite complex and sometimes unconventional. Although LAKE-RED (CVPR2024LAKERED) retrieves background embeddings from a pretrained codebook, such representations learned from generic image distributions lack camouflage-specific priors. We therefore introduce a texture-oriented background retrieval (TOBR) module that uses camouflaged images as a domain-specific background knowledge base. As illustrated in Algorithm 1, given a target foreground object, a CLIP image encoder is adopted to extract texture-aware features for the object and all candidates in the knowledge base, with the top-$k$ most similar images in the embedding space are retrieved and then incorporated into multi-modal prompt construction to guide the camouflage generation process.
3.5 Textual-Visual Condition Generator
Textual conditioning is central to T2I generation. Although CVPR2025CamoAny enhances textual prompts via prompt engineering, the incorporated background semantics remain limited. While detailed scene descriptions have been shown to improve the realism of camouflaged image synthesis (AAAI2026CTCIG), acquiring scene-specific fine-grained descriptions is often labor-intensive. We therefore propose a textual–visual condition generator (TVCG) that constructs a multi-modal prompt by combining a unified fine-grained task description with high-quality background visual embeddings. This design offers richer guidance for generation and improves visual similarity between the foreground target and the background. The multi-modal prompt comprises three components: (i) fine-grained textual task description, (ii) object category information, and (iii) reference background visual cues.
Fine-Grained Textual Task Description. Instead of using overly simplistic prompts (e.g., “image of a fish Camouflaged with Background” (CVPR2025CamoAny)) or labor-intensive, scene-specific descriptions (e.g., “A fish with mottled green and brown patterns and elongated appendages in a coral reef ecosystem.” (AAAI2026CTCIG)), we adopt a unified fine-grained textual task description “A realistic image of an object blending into its surroundings, where the background shares similar colors, textures, and patterns with the object, making it hard to distinguish. Natural lighting, photorealistic, seamless camouflage, high detail.” as the textual guidance in generation. This unified description improves generation quality while eliminating the need for manually crafting scene-specific prompts, and is encoded into textual tokens $c_{tex}$ via a CLIP text encoder.
Object Category Information. For the given image, we adopt the [CLS] token processed by a CLIP image encoder as a compact category representation of the foreground object, avoiding explicit class identification while preserving informative semantic guidance.
Algorithm 1 Texture-Oriented Background Retrieval
Input: image $\mathrm{I}$, mask $\mathrm{M}$, knowledge base with $K$ candidates $\mathrm{B}=\{S_{i}\}_{i=1}^{K}$, number of retrieval samples $k$
Initialize: similarity scores $\mathrm{S}=\varnothing$, retrieval samples $\mathrm{R}=\varnothing$, target $E_{t}=MaskedAvgPool(CLIP_{I}(\mathrm{I}),\mathrm{M})$
for $i=1$ to $K$ do
  sample $E_{i}=GlobalAvgPool(CLIP_{I}(S_{i}))$
  score $R_{i}=CosineSimilarity(E_{t},E_{i})$, $\mathrm{S}\leftarrow R_{i}$
end for
for $j=1$ to $k$ do
  index $x=ArgMax(\mathrm{S})$, $\mathrm{R}\leftarrow S_{x}$
  remove $S_{x}$ from $\mathrm{S}$
end for
Return: retrieval samples $\mathrm{R}=\{S_{j}\}_{j=1}^{k}$
Reference Background Visual Cues. To provide richer background guidance for the camouflage generation process, we incorporate the samples retrieved by TOBR as visual reference knowledge into the multi-modal prompt. Specifically, we employ a CLIP image encoder to extract visual embeddings from the input image $\mathrm{I}$ and its top-$k$ TOBR-retrieved samples $\{S_{j}\}_{j=1}^{k}$. Guided by the foreground mask $\mathrm{M}$, the foreground region of the input image is preserved, while the background regions are fused with the retrieved visual knowledge to construct background-aware visual cues $c_{vis}$. This process can be formulated as:
$e_{t}=CLIP_{I}(\mathrm{I}),\ \ e_{j}=CLIP_{I}(S_{j}),\ j=1,2,...,k,$ (4)
$c_{vis}=\left\{\begin{aligned} &e_{t}\cdot\mathrm{M}+\frac{1}{k+1}\cdot(e_{t}+\sum_{j=1}^{k}e_{j})\cdot(\mathbf{1}-\mathrm{M}),\ if\ training\\ &e_{t}\cdot\mathrm{M}+\frac{1}{k}\cdot\sum_{j=1}^{k}e_{j}\cdot(\mathbf{1}-\mathrm{M}),\ if\ inference.\end{aligned}\right$. (5)
The final multi-modal conditional prompt is the concatenation of $_{txt}, cc$$_{cls}$, and $c_{vis}$ along the sequence dimension.
Figure 4: Qualitative comparison with existing SOTA methods including style transfer based (column 3 to 4), out-painting based (column 5 to 9), and text-driven (column 10) approaches. Our results are demonstrated in the rightmost column with red box.
Table 1: Quantitative comparison with SOTA methods on LAKE-RED datasets. The 1st, 2nd and 3rd best results are highlighted in red, green and blue. $\uparrow$ / $\downarrow$ represents the higher / lower the better. ST: Style transfer based. OP: Out-paining based. TD: Text-driven.
Method Type Camouflaged Objects (6,473) Salient Objects (6,473) General Objects (6,473) Overall (19,419)
[…]
Original 1.0027 2.4837 1.6818 1.7227
AB $_{\textit{IPOL 2016}}$ ST 117.11 0.0645 0.2586 0.9200 126.78 0.0614 0.2753 1.2377 133.89 0.0645 0.2884 1.1668 120.21 0.0623 0.2741 1.1082
CI $_{\textit{TOG 2010}}$ ST 124.49 0.0662 0.2133 2.0678 136.30 0.7380 0.2117 2.2911 137.19 0.0713 0.2228 2.4039 128.51 0.0693 0.2159 2.2543
AdaIN $_{\textit{Iccv 2017}}$ ST 125.16 0.0721 0.1754 0.8821 133.20 0.0702 0.1880 1.3065 136.93 0.0714 0.2077 1.1534 126.94 0.0703 0.1904 1.1140
DCI $_{\textit{AAAI 2010}}$ ST 130.21 0.0689 0.1748 6.8849 134.92 0.0665 0.1794 5.8858 137.99 0.0690 0.2010 6.1119 130.52 0.0673 0.1850 6.2935
LCGNet $_{\textit{TMM 2023}}$ ST 129.80 0.0504 0.1760 1.6329 136.24 0.0597 0.1781 1.7395 132.64 0.0548 0.1995 1.9137 129.88 0.0550 0.1845 1.7621
TFill $_{\textit{CVPR 2022}}$ OP 63.74 0.0336 0.2527 0.3399 96.91 0.0453 0.2886 0.6743 122.44 0.0747 0.2989 0.7585 80.39 0.0438 0.2801 0.5909
LDM $_{\textit{CVPR 2022}}$ OP 58.65 0.0380 0.2390 1.0277 107.38 0.0524 0.3287 1.7082 129.04 0.0748 0.3090 1.4339 84.48 0.0488 0.2923 1.3899
RePaint-L $_{\textit{CVPR 2022}}$ OP 76.80 0.0459 0.2597 3.6712 114.96 0.0497 0.3504 4.1944 136.18 0.0686 0.3303 3.5917 96.14 0.0498 0.3135 3.8191
LAKE-RED $_{\textit{CVPR 2024}}$ OP 39.55 0.0212 0.2208 0.9191 88.70 0.0428 0.2881 1.4268 102.67 0.0555 0.2736 1.4216 64.27 0.0355 0.2608 1.2558
FACIG $_{\textit{ICME 2025}}$ OP 27.61 0.0099 82.23 0.0326 96.94 0.0503 52.87 0.0229
CamoAny $_{\textit{CVPR 2025}}$ OP 32.05 0.0140 0.1591 0.2948 52.24 0.0256 0.2052 0.5905 86.35 0.0430 0.1953 0.6173 42.85 0.0233 0.1865 0.5009
CT-CIG $_{\textit{AAAI 2026}}$ TD 30.59 0.0085 0.1972 1.0199 81.60 0.0230 0.1910 1.1938 104.46 0.0241 0.2091 1.4424 52.88 0.0169 0.1991 1.2187
RealCamo (Ours) OP 9.06 0.0015 0.3927 0.4648 11.16 0.0028 0.4670 0.9238 16.11 0.0056 0.4285 0.8365 6.93 0.0025 0.4294 0.7417
3.6 Training Objective
In RealCamo, the Stable Diffusion (SD) and the CLIP text and image encoders are all frozen, leaving ControlNet as the only trainable component. The training objective follows Equation 2, where the original control $c^{f}$ is replaced by layout controls $\{c_{\mathrm{cst}},c_{\mathrm{dep}},c_{\mathrm{hed}}\}$, and the textual condition $c^{t}$ is replaced by multi-modal conditions $\{c_{\mathrm{txt}},c_{\mathrm{cls}},c_{\mathrm{vis}}\}$. This design enables effective controllable camouflage synthesis while preserving the generalization of the pretrained SD.
3.7 Background-Foreground Distribution Divergence
To more effectively quantify the degree of camouflage in synthesized images, we propose a background–foreground distribution divergence metric, denoted as $\mathrm{KL}_{BF}$, based on KL divergence. It measures visual similarity between the generated background and the foreground target via their pixel-value distributions, where lower values indicate stronger camouflage effectiveness. Specifically, given an image $\mathrm{I}$ and its foreground mask $\mathrm{M}$, we compute the KL divergence between background and foreground distributions for each RGB channel and average them to obtain $\mathrm{KL}_{BF}$. Formally, the $\mathrm{KL}_{BF}$ can be calculated as:
$\mathrm{KL}_{c}$ $=KL(\mathrm{I}_{c}\cdot(\mathbf{1}-\mathrm{M}),\mathrm{I}_{c}\cdot\mathrm{M}),\ c\in\{R,G,B\},$ (6)
$\mathrm{KL}_{BF}$ $=Avg(\mathrm{KL}_{R},\mathrm{KL}_{G},\mathrm{KL}_{B}).$ (7)
Figure 5: Qualitative results of ablation study on main components of our proposed RealCamo framework.
Table 2: Quantitative results of ablation study on main components. The best results are highlighted in bold. $\uparrow$ / $\downarrow$ represents the higher / lower the better. For layout controls, $\times$ means only a foreground target with white background is used, without depth and HED controls. For textual-visual condition, $\times$ means only simple textual prompt is used, without additional visual tokens.
Configuration Camouflaged Objects (2,000) Salient Objects (2,000) General Objects (2,000) Overall (6,000)
[…]
$\times$ $\times$ 46.52 0.0149 0.1446 0.7371 73.38 0.0266 0.2002 1.5530 102.35 0.0340 0.1953 1.3155 46.21 0.0212 0.1800 1.2019
$\surd$ $\times$ 37.59 0.0121 0.2910 0.4019 43.97 0.0151 0.3525 0.9349 61.92 0.0280 0.3160 0.8637 27.71 0.0162 0.3198 0.7335
$\times$ $\surd$ 32.98 0.0099 0.1712 0.7405 55.01 0.0137 0.2320 1.5997 87.27 0.0237 0.2186 1.3464 33.14 0.0106 0.2073 1.2289
$\surd$ $\surd$ 31.77 0.0071 0.3254 0.3398 35.61 0.0083 0.4027 0.8659 44.90 0.0143 0.3695 0.7938 20.30 0.0083 0.3659 0.6665
Table 3: Quantitative results of ablations on layout controls. The best results are highlighted in bold. $\uparrow$ / $\downarrow$ represents the higher / lower the better.
Configuration Overall (6,000)
[…]
$\surd$ $\times$ $\times$ 32.99 0.0118 0.1917 0.6343
$\surd$ $\surd$ $\times$ 25.30 0.0104 0.2887 0.6994
$\surd$ $\times$ $\surd$ 25.54 0.0102 0.2949 0.6957
$\surd$ $\surd$ $\surd$ 20.30 0.0083 0.3659 0.6665
4 Experiments
4.1 Experimental Settings
Datasets and Metrics. We conduct experiments on the LAKE-RED dataset (CVPR2024LAKERED), which contains 4,040 training images and 19,419 images for evaluation, including 6,473 camouflaged, salient, and general images, respectively. Following CVPR2024LAKERED, we adopt FID (ICLR2018FID) and KID (NIPSl2017KID) to assess the overall quality of generated images. In addition, SSIM (TIP2004SSIM) is employed to measure structural consistency between generated images and their original references, while our proposed $\mathrm{KL}_{BF}$ metric is used to quantify the effectiveness of visual camouflage.
Implementation Details. The proposed framework is built upon Stable Diffusion 1.5 (CVPR2022LDM) augmented with ControlNet (ICCV2023ControlNet). During training, all input images, control conditions, and corresponding masks are resized to $512\times 512$, and a batch size of 4 is used. We set the control scale to 1.0 and train the model for 600,000 iterations with a learning rate of 1e-5. The generated images are also in a resolution of $512\times 512$ and are resized to their original size for evaluation. All experiments are conducted on a single NVIDIA L20 GPU.
Table 4: Quantitative results of ablations on textual-visual condition. The best results are highlighted in bold. $\uparrow$ / $\downarrow$ represents the higher / lower the better. Simple: The simple textual prompt. Detailed: The unified fine-grained textual task description.
Configuration Overall (6,000)
[…]
Simple $\times$ 46.21 0.0212 0.1800 1.2019
Detailed $\times$ 27.71 0.0162 0.3198 0.7335
$\times$ $\surd$ 33.14 0.0106 0.2073 1.2289
Detailed $\surd$ 20.30 0.0083 0.3659 0.6665
4.2 Comparison with the State-of-the-arts (SOTAs)
We compare our RealCamo with 12 SOTAs, including five style transfer based methods (AB (IPOL2016AB), CI (TOG2010Chu), AdaIN (ICCV2017AdaIN), DCI (AAAI2020DCI), and LCGNet (TMM2022LCGNet)), six out-painting based methods (TFill (CVPR2022TFill), LDM (CVPR2022LDM), RePaint-L (CVPR2022Repaint), LAKE-RED (CVPR2024LAKERED), FACIG (ICME2025FACIG), and CamoAny (CVPR2025CamoAny)), and one recently text-driven method (CT-CIG (AAAI2026CTCIG)). We re-evaluate the benchmark results provided by LAKE-RED and CT-CIG. Due to a lack of available resources, we reproduce CamoAny with the same settings as ours, and directly copy the results from the original paper for FACIG.
Table 5: Performance of MVANet (CVPR2024MVANet) on four COD benchmark datasets with different training data. The best results are highlighted in bold. $\uparrow$ / $\downarrow$ represents the higher / lower the better. Original: The traditional COD training data. Gen SOD&GOD: Synthesis images by RealCamo of the salient and general subsets in LAKE-RED evaluation set. Gen COD: Synthesis images by RealCamo of the LAKE-RED training set (i.e., our constructed SynCOD12K dataset).
Training Data CHAMELEON (76) CAMO (250) COD10K (2,026) NC4K (4,121)
$S_{\alpha}\uparrow$ $E_{\phi}\uparrow$ $F_{\beta}\uparrow$ $F_{\beta}^{w}\uparrow$ $M\downarrow$ $S_{\alpha}\uparrow$ $E_{\phi}\uparrow$ $F_{\beta}\uparrow$ $F_{\beta}^{w}\uparrow$ $M\downarrow$ $S_{\alpha}\uparrow$ $E_{\phi}\uparrow$ $F_{\beta}\uparrow$ $F_{\beta}^{w}\uparrow$ $M\downarrow$ $S_{\alpha}\uparrow$ $E_{\phi}\uparrow$ $F_{\beta}\uparrow$ $F_{\beta}^{w}\uparrow$ $M\downarrow$
Original 0.915 0.929 0.871 0.846 0.022 0.870 0.896 0.842 0.809 0.055 0.882 0.917 0.828 0.792 0.025 0.886 0.910 0.855 0.820 0.038
Gen SOD&GOD 0.905 0.927 0.846 0.796 0.031 0.826 0.842 0.760 0.717 0.079 0.865 0.896 0.784 0.719 0.035 0.879 0.897 0.821 0.769 0.048
Gen COD 0.924 0.949 0.889 0.871 0.019 0.882 0.916 0.856 0.826 0.049 0.896 0.934 0.850 0.818 0.021 0.891 0.919 0.860 0.824 0.037
Qualitative Comparison. Fig. 4 shows qualitative comparison between our proposed RealCamo and other methods on camouflaged, salient, and generic objects. The style transfer based LCGNet embeds targets into given backgrounds, achieving strong visual camouflage but severely altering target semantics. In contrast, out-painting based methods often produce insufficient camouflage or cluttered, semantically implausible backgrounds, degrading realism. Although recent text-driven CT-CIG generates visually convincing results, they fail to preserve original targets and thus require re-annotation, limiting their usefulness for downstream COD tasks. In comparison, our RealCamo achieves both high visual similarity and strong foreground–background semantic consistency while preserving target integrity, resulting in more realistic and practically valuable camouflaged images.
Quantitative Comparison. Quantitative results are summarized in Tab. 1. Compared to early style transfer based methods, latest out-painting based and text-driven approaches achieve obviously better image quality, indicated by lower FID and KID. Owing to the introduction of explicit layout controls and textual–visual multi-modal conditioning, our RealCamo attains highly competitive generation performance. Although it does not achieve the lowest $\mathrm{KL}_{BF}$ score, our method balances visual camouflage with structural and semantic fidelity, producing more realistic images. By revisiting Fig. 4, methods with lower $\mathrm{KL}_{BF}$ values (e.g., CamoAny and TFill) indeed yield stronger visual camouflage, further validating our proposed $\mathrm{KL}_{BF}$ as an effective metric for quantifying camouflage effectiveness.
User Study. We further conduct a user study to comprehensively evaluate the quality of synthesized camouflaged images based on human feedback. Detailed experimental settings, evaluation criteria, and statistical results and analyses will be reported in Appendix C.
4.3 Ablation Study
We conduct ablation studies to validate the proposed components. Due to computational constraints, all variants are trained for 100,000 iterations, and evaluation is performed on 6,000 images uniformly sampled from three test subsets. Results in Tab. 2 and Fig. 5 show that explicit layout controls improve structural consistency and overall image quality, supporting our claim that “the semantics of images are largely determined by their structural layout”. Incorporating textual–visual condition further enhances visual realism, as well as foreground–background visual similarity and semantic consistency. Additional results of ablations on individual layout control and textual–visual components are reported in Tabs 3 and 4. More quantitative and qualitative result with further analyses will be provided in Appendix.
4.4 Downstream Application
We further evaluate the impact of synthesized camouflaged images on downstream COD application. An awesome dichotomous segmentation model, MVANet (CVPR2024MVANet), is trained with different data configurations and evaluated on four COD benchmarks. Quantitative results and prediction results are presented in Tab. 5 and Fig. 6. We can observe that, converting salient or general images into camouflaged ones degrades COD performance, possibly due to domain gaps between these synthesized samples and real COD data. In contrast, training with camouflaged-synthesis images consistently improves COD performance, highlighting the effectiveness of our generation approach for data augmentation. We also examine the impact of synthesized images on salient and general object detection. More details of experimental settings, metrics and results, and the constructed SynCOD12K dataset, will be reported in Appendix.
Figure 6: Prediction results with different training data.
5 Conclusion
In this paper, we propose RealCamo, a unified framework for controllable camouflaged image generation. RealCamo explicitly incorporates additional layout controls to regulate the structural composition of generated images, thereby enhancing semantic consistency between foreground objects and synthesized backgrounds. In addition, it constructs a multi-modal textual–visual condition by combining unified fine-grained textual task descriptions with texture-oriented background retrieved visual cues , which jointly guides the generation process to improve visual quality and realism. Extensive experiments and qualitative analyses validate the effectiveness of the proposed framework and demonstrate its strong potential for enhancing downstream camouflaged object detection applications.
Appendix A Camouflage Object Detection
A.1 Related Work
As the extension of general object detection (GOD) (CVPR2015FCN-GOD) and salient object detection (SOD) (TIP2015SOD), camouflaged object detection (COD) (CVPR2020SINet-COD10K; ICML2025RUN) is proposed to accurately localize and precisely segment camouflaged targets hidden within complex scenes, which is applicable to numerous real-world scenarios, such as concealed defect detection in industry, pest monitoring in agriculture, and medical lesions diagnosis (AIR2024CODsurvey). SINet (CVPR2020SINet-COD10K) pioneered COD using a biologically inspired two-stage search-and-identify framework, followed by a wide range of bio-inspired (CVPR2021PFNet; TPAMI2022SINetV2; CVPR2022ZoomNet; CVPR2022SegMaR; TIP2023FSNet; TIP2024IdeNet) and feature-enhanced (CVPR2021MGL; CVPR2021LSR-NC4K; TIP2022FAPNet; CVPR2023FSPNet; ICLR2024ICEG; TPAMI2024CamoFormer) methods. Subsequent studies incorporated auxiliary cues to complement RGB images, including edges and textures (IJCAI2022BgNet; MIR2023DGNet; CVPR2023FEDER), depth information (ICCV2023PopNet; ACMMM2023DaCOD; ACMMM2024DSAM), and frequency-domain representations (ACMMM2023FRINet; ACMMM2023FPNet; ECCV2024FSEL). Representative works exploit frequency decomposition with boundary refinement (CVPR2023FEDER), depth-induced saliency enhancement (ICCV2023PopNet), and joint frequency–spatial learning (ECCV2024FSEL). Despite steady progress, the development of COD remains fundamentally limited by the small scale of available training datasets, which drives the development of camouflaged image generation tasks.
A.2 Datasets and Evaluation Metrics
Datasets. We evaluate the effectiveness of our synthetic dataset, SynCOD12K, on the COD task following the standard practice of SINet (CVPR2020SINet-COD10K). Experiments are conducted on four widely used benchmark datasets: CHAMELEON (CHAMELEON), CAMO (CVIU2019CAMO), COD10K (CVPR2020SINet-COD10K), and NC4K (CVPR2021LSR-NC4K). The CHAMELEON dataset contains 76 images, while CAMO consists of 1,250 images across 8 classes. COD10K includes 5,066 images spanning ten super-classes, and NC4K serves as the largest testing benchmark with 4,121 well-annotated images. For standard COD training, we follow the common practice by using 1,000 images from CAMO and 3,040 images from COD10K as the training set, which also aligns with the training protocol for camouflaged image generation in (CVPR2024LAKERED). The remaining images from these two datasets, together with all images from CHAMELEON and NC4K, are used for evaluation.
Evaluation Metrics. To thoroughly evaluate the performance, we adopt five standard metrics widely used in COD and SOD tasks: mean absolute error ($M$) (MAE), mean E-measure ($E_{\phi}$) (E-measure), mean F-measure ($F_{\beta}$) (F-measure), weighted F-measure ($F_{\beta}^{w}$) (wF-measure), and structure measure ($S_{\alpha}$) (S-measure). Better performance is indicated by lower values of $M$ and higher values of $E_{\phi}$, $F_{\beta}$, $F_{\beta}^{w}$, and $S_{\alpha}$.
Appendix B Preliminary
B.1 Diffusion Models
Tex-to-image (T2I) generation methods, represented by Stable Diffusion (CVPR2022LDM), synthesize images conditioned on textual descriptions ($c^{t}$) via a denoising diffusion paradigm (NIPS2020DDPM). These models can operate either in pixel space or latent space, with latent diffusion models (CVPR2022LDM) offering substantially improved computational efficiency.
Forward Process. In diffusion models, noise sampled from the Gaussian distribution $\epsilon\sim\mathcal{N}(0,I)$ is progressively added to a clean image $x_{0}$, resulting in a noisy sample $x_{t}$ at timestep $t$:
$x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,$ (8)
where $\alpha_{t}=1-\beta_{t}$, $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$, and $\{\beta_{t}\}$ follows a variance schedule.
Training Objective. A denoising network $\epsilon_{\theta}$ is trained to predict the injected noise $\epsilon$ by minimizing the following objective:
$\mathcal{L}=\mathbb{E}_{x_{0},t,\epsilon\sim\mathcal{N}(0,I)}\left[\parallel\epsilon-\epsilon_{\theta}(x_{t},t,c^{t})\parallel_{2}^{2}\right].$ (9)
Reverse Process. After training the denoising network, an image $x_{0}$ can be gradually obtained by repeating the denoise operation $T$ times on an entire Gaussian noise $x_{T}\sim\mathcal{N}(0,I)$, following:
$x_{t-1}=\frac{1}{\sqrt{\bar{\alpha}_{t}}}\left(x_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(x_{t},t,c^{t})\right)+\sigma_{t}\epsilon^{\prime},$ (10)
where $t$ starts from $T$, $\epsilon^{\prime}\sim\mathcal{N}(0.I)$, and $\sigma_{t}^{2}=\beta_{t}$. Intermediate estimations of $x_{0}$ can also be directly obtained at any timestep $t$ following:
$\hat{x}_{0}=\frac{x_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}(x_{t},t,c^{t})}{\sqrt{\bar{\alpha}_{t}}}.$ (11)
In latent diffusion models (LDMs) (CVPR2022LDM), the above forward and reverse process is conducted on latent embeddings $z=\mathcal{E}_{\mathrm{VAE}}(x)$, as a latent representation of raw RGB pixels $x$ encoded by a VAE (ICLR2014VAE1; ICML2014VAE2). And the denoised latent representation $z_{0}$ is further decoded into a normal RGB image using the VAE decoder, as $x_{0}=\mathcal{D}_{\mathrm{VAE}}(z_{0})$.
B.2 ControlNet
ControlNet (ICCV2023ControlNet) remarkably extends the controllability of large pretrained T2I LDMs by incorporating addition conditions ($c^{f}$) like Canny edges, human pose, or depth map, which provide structural priors for the target generated image.
Architecture. ControlNet controls the generating process by injecting additional conditions into the blocks of a pretrained neural network. With the definition $\mathcal{F}(\cdot;\Theta)$ of a trained neural block with parameters $\Theta$, that transform an input feature map $x$ into an output feature map $y$ as $y=\mathcal{F}(x;\Theta)$, a trainable copy of $\mathcal{F}$ is created as $\mathcal{F}_{c}(\cdot;\Theta_{\mathrm{c}})$.
This trainable copy takes an external conditioning vector $c^{f}$ as input, and is connected to the locked model $\mathcal{F}$ with zero convolution layers $\mathcal{Z}(\cdot;\cdot)$, which are $1\times 1$ convolution layers with both weight and bias initialized to zeros. The ControlNet then generated the output $y_{\mathrm{c}}$ by two instances of zero convolutions with parameters $\Theta_{\mathrm{Z}1}$ and $\Theta_{\mathrm{Z}2}$ as:
$y_{\mathrm{c}}=\mathcal{F}(x;\Theta)+\mathcal{Z}(\mathcal{F}(x+\mathcal{Z}(c^{f};\Theta_{\mathrm{Z}1});\Theta_{\mathrm{c}});\Theta_{\mathrm{Z}2})$ (12)
Training Objective. With the introduction of additional conditioning vector $c^{f}$, the overall training objective (Equation 9) is reformulated as:
$\mathcal{L}=\mathbb{E}_{x_{0},t,c^{t},c^{f},\epsilon\sim\mathcal{N}(0.I)}\left[\parallel\epsilon-\epsilon_{\theta}(x_{t},t,c^{t},c^{f})\parallel_{2}^{2}\right].$ (13)
This enables fine-grained spatial control during generation, allowing precise alignment between generated content and structural constraints.
Appendix C User Study
The results are currently being collected and pending further analysis..